The position of a squirel running in a park s gven The position of a squirrel running in a park is given by (0.280 m/s)t + (0.0360 m/t) + (0.0190 m/s3)tj

D) At 5.24 s , what is the magnitude of the squirrel's velocity?

E) At 5.24 s , what is the direction (in degrees counterclockwise from +x-axis) of the squirrel's velocity?
Answer #1

D) v = dr/dt

= ( 0.28 + 0.036*2*t)i + 0.019*3*t^2 j

at t = 5.24 s

v = (0.28 + 0.036*2*5.4) i + 0.019*3*5.24^2 j

= 0.6688 i + 1.565 j

|v| = sqrt(0.6688^2 + 1.565^2)

= 1.7 m/s

E) direction : theta = tan^-1(vy/vx)

= tan^-1(1.565/0.6688)

= 66.86 degrees